# Multiscale Hodge Scattering Networks for Data Analysis

### v2

@@ 3 Multiscale Hodge Scattering Transform
Let the $\kappa$-HGLET or $\kappa$-GHWT dictionary vectors be arranged as ${\boldsymbol{\Phi}}^{J}\,:=\,\left\{\Phi^{j}\right\}_{j=0}^{J}$ where each $\Phi^{j}$ is an ONB at scale (or level) $j$ with $j=0$ being the finest scale basis, composed of delta functions. Note that this definition of the scale parameter $j$ is the opposite of that used in the previous sections. In general, we have ${j_{\mathrm{max}}}+1\approx\log_{2}n+1$ different levels given by the hierarchical bipartition scheme, but in practice, the features extracted by large $j$ values are not very descriptive gao2019geometric . Hence, we typically use the first $J+1(\leq{j_{\mathrm{max}}}+1)$ levels.
@@
where $0\leq j^{(1)}<\cdots<j^{(m)}\leq J$. However, due to the combinatorial blow-up in the number of features at each order, it is rare to use more than 2nd or 3rd-order features. Note that the $m$th order features are computed by applying a multiplication with the appropriate (sparse) weight $\Phi^{j}$ then applying a pointwise nonlinearity to $(m-1)^{\text{st}}$-order features. Further, as we later find in our numerical experiments, high-order moments $(Q>4)$ are not very useful in practice due to their instability BRUNA-MALLAT ; chew2022geometric ; blank. We notate the transform without or the sum or normalization actor as $\hat{S}$. We can write the higher order features as:
@@
$\cdots=S^{1}\left(q,j^{(m)})\hat{S}^{1}(1,j^{(m-1)}\right)\cdots\hat{S}^{1}\left(1,j^{(1)}\right){\boldsymbol{f}}$
This behavior mimics the architecture of convolutional neural networks (with fixed weights) and has been studied extensively MALLAT-SCAT . However, unlike traditional feed-forward networks, rather than only considering the features extracted at end of the cascade of linear and nonlinear layers, we concatenate the representation extracted at each layer for downstream analysis. In general, we refer to the process of extracting these features as the Multiscale Hodge Scattering Transform (MHST). Later we will analyze these features with relatively simple machine learning models such as support vector machines (SVMs) and logistic regression models (LRMs). When working with such models with learnable parameters in concert with these deterministic features, we refer to it as the Multiscale Hodge Scattering Network (MHSN).
@@ 3.1 Local Pooling
In general, we can gather all of the moments $\leq Q$ and of orders $\leq M$ to have a total of $Q\sum_{k=0}^{M}\binom{J+1}{k}$ features for a given signal. The summations from $i=1$ to $i=n$ in (3)–(5) can be viewed as global pooling operations. In situations where permutation invariance is not required (i.e., all signals are defined on a fixed complex with known node ordering), however, we can omit these sums, i.e., no pooling is done. As a result, we are left with $Qn\sum_{k=0}^{M}\binom{J+1}{k}$ features for each signal. We can also generate intermediate features between these two extremes: retain sums over each region $k$ at level $j$ instead of not summing at all or summing all the regions of level $j$ in (3)–(5). This can be viewed as local pooling operations and results in a tuple of features rather than a single value as in the original scattering transform. This is similar to the windowed scattering transforms recently proposed by chew2022geometric , but here we leverage the multiscale decomposition provided by the hierarchical partition tree to avoid introducing a new user parameter to determine the window size. In the case of these local pools, we replace the normalization factor in the averaging operation with the number of elements in the local pool rather than the number in the entire simplex as defined in Eq. (7). This gives us a total of $Q\sum_{j=0}^{M}K^{j}\binom{J+1}{j}$ features, where $K^{j}$ is the number of local sums taken in the $j$th level, i.e., the number of regions at scale $j$.
We denote these transforms as $S^{m}_{j}$ where $j$ denotes the level at which the sum has taken place. So $S^{m}_{j_{\mathrm{max}}}=S^{m}$ indicated the standard max-pooling-like scheme as defined by BRUNA-MALLAT ; gao2019geometric . Then $S^{m}_{0}=\hat{S}$ denotes the transform without any sums and $S^{m}_{2}$ denotes the transform with local pools determined by the third level (i.e., $j=2$) of the partition tree. In general we have:
@@ 4 Theoretical Analysis of Multiscale Hodge Scattering Transforms
where $\alpha$ is a constant in $(0,1]$. With these definitions, the dictionary coefficient decay and approximation results of sharon2015class ; IRION-SAITO-TSIPN for the GHWT and HGLET can be applied to the $\kappa$-GHWT and $\kappa$-HGLET bases as detailed further in SAITO-SCHONSHECK-SHVARTS .
@@
See Theorem 3.2 of IRION-SAITO-TSIPN and Theorem 6.3 of sharon2015class . ∎
@@
This remark is clearly true since ${\mathscr{W}}$ is simply a collection of $J+1$ orthogonal matrices, and orthogonal transforms preserve the $\ell^{2}$-norm. Although trivial, this fact will be vital for later proofs. Next we show that the MHST is a non-expansive operation. This allows for powerful nonlinear analysis properties lohmiller1998contraction which we detail later.
@@
Next we show that our networks are invariant under group operations which preserve the inner product $\langle\cdot,\cdot,\rangle_{w}$, such as permutations of the $\kappa$-simplices. That is, relabeling indices of the elements of $C_{\kappa}$ does not affect the output of the transform. For example, given a signal ${\boldsymbol{f}}$ defined on the triangular faces of a simplex, the indexing of the triangles does not effect the globally pooled transform and permuting this indexing results in an equivalent permutation of the non-pooled signal. This theorem is analogous to Theorem 3 in chew2022geometric and Proposition 4.1 in zou2020graph , but generalizes to any $\kappa$, rather than strictly to the nodes ($\kappa$=0 case).
@@ 5 Signal Classification
We first demonstrate the effectiveness of our MHSNs with the article category classification problem using the Science News database SCIENCE-NEWS ; IRION-SAITO-MLSP16 . We do not claim state-of-the-art results for this problem, but rather use it to illustrate the advantages of analyzing signals via higher-dimensional simplices. We also show how locally-pooled networks can shatter problems in which traditional globally-pooled scattering networks fail to differentiate between classes. The Science News dataset contains 1042 scientific news articles classified into eight fields: Anthropology, Astronomy; Behavioral Sciences; Earth Sciences; Life Sciences; Math/CS; Medicine; Physics. Each article is tagged with keywords from a pool of 1133 words selected by the database curators. We determine a simplicial complex from these keywords by computing their word2vec rong2014word2vec embeddings based on Google’s publicly available pre-trained model google-word2vec-site . We generate a symmetric $k$-nearest neighbor graph of the embedded words and then generate $\kappa$-simplices of the graph. Therefore, a $\kappa$-simplex in this keyword graph corresponds to $\kappa$-face, which represents a combination of $\kappa+1$ words.
@@
Tables 1 and 2 compare the performance of our proposed methods with the other simpler methods, i.e., the raw expansion coefficients of the signals relative to the standard ONBs (Delta; Fourier) or the dictionaries (Diffusion; HGLET; GHWT). The parameters for the feature computations were set as $(J,M,Q)=(5,3,4)$. For each $\kappa$, we performed the five-fold cross-validation, i.e., we randomly split these 1042 signals into 80% training and 20% test sets and repeat the experiment 5 times with different train/test splits. In every case we use the $\ell^{2}$-regularized LRM provided by scikit-learn scikit-learn without any additional hyperparameter tuning to compute the classification.
@@
We also observe that the networks based on $\kappa$-HGLET and $\kappa$-GHWT generally outperform those based on Diffusion Wavelets. This is likely due to the highly localized and piecewise constant nature of the input signals, which are well-approximated by these dictionaries SAITO-SCHONSHECK-SHVARTS . In the next section, where the signals are not localized, we do not observe this difference.
Table 1: Article category classification accuracy for $5$-NN graph of the Science News dataset for different simplex degrees. Dict. implies that the SVM is trained solely on the dictionary coefficients while GP, LP, NP imply scattering networks with global, local, and no pooling, respectively
@@
. The best performer for each $\kappa$ is indicated in bold while the underlined bold numbers are the best among all $\kappa$’s.
Table 1: Article category classification accuracy for $5$-NN graph of the Science News dataset for different simplex degrees. Dict. implies that the SVM is trained solely on the dictionary coefficients while GP, LP, NP imply scattering networks with global, local, and no pooling, respectively
@@ 6 Domain Classification
Another vital application of geometric scattering networks and graph neural networks (GNNs) is graph (and simplex) classification. Broadly speaking, this problem consists of predicting a label of a social or chemical graph based on a training set of similar graphs with different configurations (i.e., different numbers of nodes and edges). Since the size of the graphs varies greatly within these datasets, we employ only the global-pooling version of our MHSN, akin to the previous efforts reported in gao2019geometric ; chew2022geometric , which were based on geometric scattering methods. We employ a simple SVM with Gaussian radial basis functions to classify the features extracted by the MHSTs. Moreover, we tune the hyperparameters controlling the strength of the $\ell^{2}$-regularization and the kernel coefficients via the cross-validation scheme presented in gao2019geometric using the same search space and validation indexes.
In some applications, there is additional labeling (node labels, edge labels, etc.), but we consider the unlabeled case for now. Therefore, we have to compute permutation-invariant input features based only on topological/geometric information from the boundary matrices. Since many of the graphs are too small to contain high-degree simplices, we only consider node and edge-based features and transforms. For the node-based signals, we first compute the eccentricity and clustering coefficient of each node. Then, pass each into the scattering transform and concatenate the outputs (using global pooling). Finally, we used these features as inputs to an SVM. For the edge-based signals, we use two sets of features: 1) the number of off-diagonal terms in Hodge Laplacian; and 2) the average degree of the head and tail nodes of each edge. For each problem we train three models: 1) using only node features; 2) using only edge features; and 3) using both sets. We adopt Gaussian radial basis functions with hyperparameter space and tuning scheme proposed by gao2019geometric .
We compare these results with those obtained by the geometric scattering network using SVM (GS-SVM) as well as several popular GNN models including the graph convolution network (GCN) kipf2017semi , universal graph transform (UGT) nguyen2022universal , dynamic graph CNN (DGCNN) wang2018dynamic , graph attention network (GAT) velickovic2018graph , and graph feature network (GFN) chen2019powerful . For each competing method, we reproduce the results in the significant figures reported in their original publications; we report to 2 decimal places for our method. More information on the benchmark datasets can be found in A. We remark that, as of this article’s writing, this collection of networks achieves state-of-the-art results on these datasets according to the Papers with Code Leaderboards paperwithcode . Further details on these datasets and their associated classification problems are presented in  A and the references therein.
@@ 7 Molecular Dynamics
Our MHSNs can also be used for regression problems where the goal is to predict a continuous property of a simplicial complex (or simply a graph) based on a set of observations of the complex under various conditions. Therefore, they are quite suitable for learning molecular dynamics, particularly the potential energy surface of a molecule, given a few registrations of the molecule and its energies. The Revised Molecular Dynamics 17 (rMD17 dataset) bowman2022md17 contains 100,000 structures and associated energies of various molecules. However, these structures are taken from a molecular dynamics simulation, i.e., time series data, which is not independent and identically distributed. To overcome this, instead of using the entire dataset, we use a set of five standardized 80/20 training/test indexes of 1,000 snapshots of the structures/energies per molecule selected by the authors of the dataset bowman2022md17 .
@@
We extract a rich set of features for each structure (i.e., a pose or confirmation of a molecule) using our MHSNs and then employ a support vector regression (SVR) method with Gaussian radial basis functions to approximate the energy. More specifically, for each molecule, we first compute the average position of each atom across the training set. Then, using these positions, we create a $k$NN-graph (with $k=5$) as a template and compute MHST coefficients for nodes and edges of this simplicial complex. By using this graph, rather than the molecular-bond graph, we can better capture the geometric information in the pose of the molecule as detailed in schutt2017schnet ; schutt2023schnetpack . Then, for each pose, we extract a set of features based only on the coordinates of the confirmation. For the nodes-based networks, we compute the Euclidean distance matrix of the node coordinates and then use each column as an input to the MHSN. For the edge-based networks, we use a diagonal matrix of the edge lengths instead. We do not use higher-degree simplices because the molecules are too small.
@@
Table 5 shows our results for aspirin (21 atoms) and paracetamol (20 atoms) molecules. We compare our MHSNs with several state-of-the-art GNN approaches designed specifically for processing molecular dynamics, including SchNet schutt2017schnet , PaiNN schutt2021equivariant , and two variants of special orthogonal networks batzner20223 ; schutt2023schnetpack . We report both the mean average error (MAE) and root mean square error (RMSE) of the energy prediction, which are the standard metrics in the computational chemistry literature. Our MHSNs perform competitively with these approaches while employing roughly 1% as many learnable parameters as the competing methods. Additionally, we again observe that our edge-based analysis outperforms the node-based analysis. This demonstrates that higher-dimensional simplex analysis can be more powerful than node-only approaches, even in cases where the underlying graph may not have many higher-dimensional structures
Diff+SVR HGLET+SVR GHWT+SVR SchNet PaiNN SO3Net I SO3Net II
@@
Table 5: Comparison of the performance of our MHSNs and the other state-of-the-art GNNs for nuclear energy prediction. We report the accuracy via MAE and RMSE as-well as the number of trainable parameters in each network

### v4

@@ 3 Multiscale Hodge Scattering Transform
Let the $\kappa$-HGLET or $\kappa$-GHWT dictionary vectors be arranged as ${\boldsymbol{\Phi}}^{J}\,:=\,\left\{\Phi^{j}\right\}_{j=0}^{J}$ where each $\Phi^{j}$ is an ONB at scale (or level) $j$ with $j=0$ being the finest scale basis, composed of delta functions. Note that this definition of the scale parameter $j$ is the opposite of that used in the previous sections. In general, we have ${j_{\mathrm{max}}}+1\approx\log_{2}n+1$ different levels given by the hierarchical bipartition scheme, but in practice, the features extracted by large $j$ values are not very descriptive [4]. Hence, we typically use the first $J+1(\leq{j_{\mathrm{max}}}+1)$ levels.
@@
where $0\leq j^{(1)}<\cdots<j^{(m)}\leq J$. However, due to the combinatorial blow-up in the number of features at each order, it is rare to use more than 2nd or 3rd-order features. Note that the $m$th order features are computed by applying a multiplication with the appropriate (sparse) weight $\Phi^{j}$ then applying a pointwise nonlinearity to $(m-1)^{\text{st}}$-order features. Further, as we later find in our numerical experiments, high-order moments $(Q>4)$ are not very useful in practice due to their instability [2, 4, 25]. {hide} We notate the transform without the sum or normalization factor as $\hat{S}$. We can write the higher order features as:
@@
$\cdots=S^{1}\left(q,j^{(m)}\right)\hat{S}^{1}\left(1,j^{(m-1)}\right)\cdots\hat{S}^{1}\left(1,j^{(1)}\right){\boldsymbol{f}}$
By defining the operator $\hat{S}^{1}(j,q){\boldsymbol{f}}\,:=\,\left|\Phi^{j}{\boldsymbol{f}}\right|^{q}$, which maps ${\boldsymbol{f}}\in{\mathbb{R}}^{n}$ to ${\mathbb{R}}_{\geq 0}^{n}$, we can also rewrite the higher-order features Eq. (5) in a more comprehensible manner as:
$S^{m}\left(q,j^{(1)},\ldots,j^{(m)}\right){\boldsymbol{f}}=S^{1}\left(q,j^{(m)}\right)\hat{S}^{1}\left(1,j^{(m-1)}\right)\cdots\hat{S}^{1}\left(1,j^{(1)}\right){\boldsymbol{f}}$ (7)
This behavior mimics the architecture of convolutional neural networks (with fixed weights) and has been studied extensively [1]. However, unlike traditional feed-forward networks, rather than only considering the features extracted at end of the cascade of linear and nonlinear layers, we concatenate the representation extracted at each layer for downstream analysis. In general, we refer to the process of extracting these features as the Multiscale Hodge Scattering Transform (MHST). Later we will analyze these features with relatively simple machine learning models such as support vector machines (SVMs) and logistic regression models (LRMs). When working with such models with learnable parameters in concert with these deterministic features, we refer to it as the Multiscale Hodge Scattering Network (MHSN).
@@ 3.1 Local Pooling
In general, we can gather all of the moments $\leq Q$ and of orders $\leq M$ to have a total of $Q\sum_{k=0}^{M}\binom{J+1}{k}$ features for a given signal. The summations from $i=1$ to $i=n$ in (3)–(5) can be viewed as global pooling operations. In situations where permutation invariance is not required (i.e., all signals are defined on a fixed complex with known node ordering), however, we can omit these sums, i.e., no pooling is done. As a result, we are left with $Qn\sum_{k=0}^{M}\binom{J+1}{k}$ features for each signal. We can also generate intermediate features between these two extremes: retain sums over each region $k$ at level $j$ instead of not summing at all or summing all the regions of level $j$ in (3)–(5). This can be viewed as local pooling operations and results in a tuple of features rather than a single value as in the original scattering transform. This is similar to the windowed scattering transforms recently proposed by [25], but here we leverage the multiscale decomposition provided by the hierarchical partition tree to avoid introducing a new user parameter to determine the window size. In the case of these local pools, we replace the normalization factor in the averaging operation with the number of elements in the local pool rather than the number in the entire simplex as defined in Eq. (8). This gives us a total of $Q\sum_{j=0}^{M}K^{j}\binom{J+1}{j}$ features, where $K^{j}$ is the number of local sums taken in the $j$th level, i.e., the number of regions at scale $j$.
We denote these transforms as $S^{m}_{j}$ where $j$ denotes the level at which the sum has taken place. So $S^{m}_{j_{\mathrm{max}}}=S^{m}$ indicated the standard max-pooling-like scheme as defined by [2, 4]. Then $S^{m}_{0}$ denotes the transform without any sums while $S^{m}_{2}$ denotes the transform with local pools determined by the third level (i.e., $j=2$) of the partition tree. In general we have:
@@ 4 Theoretical Analysis of Multiscale Hodge Scattering Transforms
where $\alpha$ is a constant in $(0,1]$. With these definitions, the dictionary coefficient decay and approximation results of [32, 6] for the GHWT and HGLET can be applied to the $\kappa$-GHWT and $\kappa$-HGLET bases as detailed further in [7].
@@
See Theorem 3.2 of [6] and Theorem 6.3 of [32]. ∎
@@
This remark is clearly true since ${\mathscr{W}}$ is simply a collection of $J+1$ orthogonal matrices, and orthogonal transforms preserve the $\ell^{2}$-norm. Although trivial, this fact will be vital for later proofs. Next we show that the MHST is a non-expansive operation. This allows for powerful nonlinear analysis properties [33] which we detail later.
@@
Next we show that our networks are invariant under group operations which preserve the inner product $\langle\cdot,\cdot,\rangle_{w}$, such as permutations of the $\kappa$-simplices. That is, relabeling indices of the elements of $C_{\kappa}$ does not affect the output of the transform. For example, given a signal ${\boldsymbol{f}}$ defined on the triangular faces of a simplex, the indexing of the triangles does not effect the globally pooled transform and permuting this indexing results in an equivalent permutation of the non-pooled signal. This theorem is analogous to Theorem 3 in [25] and Proposition 4.1 in [34], but generalizes to any $\kappa$, rather than strictly to the nodes ($\kappa$=0 case).
@@ 5 Signal Classification
We first demonstrate the effectiveness of our MHSNs with the article category classification problem using the Science News database [35, 36]. We do not claim state-of-the-art results for this problem, but rather use it to illustrate the advantages of analyzing signals via higher-dimensional simplices. We also show how locally-pooled networks can shatter problems in which traditional globally-pooled scattering networks fail to differentiate between classes. The Science News dataset contains 1042 scientific news articles classified into eight fields: Anthropology, Astronomy; Behavioral Sciences; Earth Sciences; Life Sciences; Math/CS; Medicine; Physics. Each article is tagged with keywords from a pool of 1133 words selected by the database curators. We determine a simplicial complex from these keywords by computing their word2vec [37] embeddings based on Google’s publicly available pre-trained model [38]. We generate a symmetric $k$-nearest neighbor graph of the embedded words and then generate $\kappa$-simplices of the graph. Therefore, a $\kappa$-simplex in this keyword graph corresponds to $\kappa$-face, which represents a combination of $\kappa+1$ words.
@@
Tables 1 and 2 compare the performance of our proposed methods with the other simpler methods, i.e., the raw expansion coefficients of the signals relative to the standard ONBs (Delta; Fourier) or the dictionaries (Diffusion; HGLET; GHWT). The parameters for the feature computations were set as $(J,M,Q)=(5,3,4)$. For each $\kappa$, we performed the five-fold cross-validation, i.e., we randomly split these 1042 signals into 80% training and 20% test sets and repeat the experiment 5 times with different train/test splits. In every case we use the $\ell^{2}$-regularized LRM provided by scikit-learn [39] without any additional hyperparameter tuning to compute the classification.
@@
We also observe that the networks based on $\kappa$-HGLET and $\kappa$-GHWT generally outperform those based on Diffusion Wavelets. This is likely due to the highly localized and piecewise constant nature of the input signals, which are well-approximated by these dictionaries [7]. In the next section, where the signals are not localized, we do not observe this difference.
@@
Table 2: Article category classification accuracy for $10$-NN graph of the Science News dataset for different simplex degrees. Dict. implies that the SVM is trained solely on the dictionary coefficients while GP, LP, NP imply scattering networks with global, local, and no pooling, respectively. The best performer for each $\kappa$ is indicated in bold while the underlined bold numbers are the best among all $\kappa$’s.
@@ 6 Domain Classification
Another vital application of geometric scattering networks and graph neural networks (GNNs) is graph (and simplex) classification. Broadly speaking, this problem consists of predicting a label of a social or chemical graph based on a training set of similar graphs with different configurations (i.e., different numbers of nodes and edges). Since the size of the graphs varies greatly within these datasets, we employ only the global-pooling version of our MHSN, akin to the previous efforts reported in [4, 25], which were based on geometric scattering methods. We employ a simple SVM with Gaussian radial basis functions to classify the features extracted by the MHSTs. Moreover, we tune the hyperparameters controlling the strength of the $\ell^{2}$-regularization and the kernel coefficients via the cross-validation scheme presented in [4] using the same search space and validation indexes.
In some applications, there is additional labeling (node labels, edge labels, etc.), but we consider the unlabeled case for now. Therefore, we have to compute permutation-invariant input features based only on topological/geometric information from the boundary matrices. Since many of the graphs are too small to contain high-degree simplices, we only consider node and edge-based features and transforms. For the node-based signals, we first compute the eccentricity and clustering coefficient of each node. Then, pass each into the scattering transform and concatenate the outputs (using global pooling). Finally, we used these features as inputs to an SVM. For the edge-based signals, we use two sets of features: 1) the number of off-diagonal terms in Hodge Laplacian; and 2) the average degree of the head and tail nodes of each edge. For each problem we train three models: 1) using only node features; 2) using only edge features; and 3) using both sets. We adopt Gaussian radial basis functions with hyperparameter space and tuning scheme proposed by [4].
We compare these results with those obtained by the geometric scattering network (with Diffusion Wavelets) using SVM (GS-SVM) as well as several popular GNN models including the graph convolution network (GCN) [40], universal graph transform (UGT) [41], dynamic graph CNN (DGCNN) [42], graph attention network (GAT) [43], and graph feature network (GFN) [44]. For each competing method, we reproduce the results in the significant figures reported in their original publications; we report to 2 decimal places for our method. More information on the benchmark datasets can be found in A. We remark that, as of this article’s writing, this collection of networks achieves state-of-the-art results on these datasets according to the Papers with Code Leaderboards [45]. Further details on these datasets and their associated classification problems are presented in  A and the references therein.
@@ 7 Molecular Dynamics
Our MHSNs can also be used for regression problems where the goal is to predict a continuous property of a simplicial complex (or simply a graph) based on a set of observations of the complex under various conditions. Therefore, they are quite suitable for learning molecular dynamics, particularly the potential energy surface of a molecule, given a few registrations of the molecule and its energies. The Revised Molecular Dynamics 17 (rMD17 dataset) [46] contains 100,000 structures and associated energies of various molecules. However, these structures are taken from a molecular dynamics simulation, i.e., time series data, which is not independent and identically distributed. To overcome this, instead of using the entire dataset, we use five sets of molecule snapshots and the associated potential energies. Each of these sets consists of 1,000 snapshot/energy pairs and is grouped into 800 training and 200 test samples selected by the authors of the dataset [46].
@@
We extract a rich set of features for each structure (i.e., a pose or conformation of a molecule) using our MHSNs and then employ a support vector regression (SVR) method with Gaussian radial basis functions to approximate and predict the energy. More specifically, for each molecule, we first compute the average position of each atom across the training set. Then, using these positions, we create a $k$NN-graph (with $k=5$) as a template simplicial complex. Note that by using this simplicial complex, rather than the molecular-bond graph, we can better capture the geometric information in the pose of the molecule as detailed in [47, 48]. We then generate the node and edge signals of both training and test sets as follows. We compute the Euclidean distance matrix of the node coordinates of each snapshot. For each node in the simplicial complex, we assign the corresponding column vector of the distance matrix as its node value. This generates a vector-valued node signal per snapshot. As for an edge signal for a snapshot, we extract edge lengths from the above distance matrix and create a diagonal matrix of edge lengths. Then, we assign the corresponding column vector of the edge distance matrix as its edge value. This generates a vector-valued edge signal per snapshot. We do not use simplices of dimension greater than $1$ because the molecules are too small.
@@
Table 5 shows our results for aspirin (21 atoms) and paracetamol (20 atoms) molecules. We compare our MHSNs with several state-of-the-art GNN approaches designed specifically for processing molecular dynamics, including SchNet [47], PaiNN [49], and two variants of Special Orthogonal Networks (SO3Nets) [50, 48]. We report both the mean absolute error (MAE) and root mean square error (RMSE) of the energy prediction, which are the standard metrics in the computational chemistry literature. Our MHSNs perform competitively with these approaches while employing roughly 1% as many learnable parameters as the competing methods. Additionally, we again observe that our edge-based analysis outperforms the node-based analysis. This demonstrates that higher-dimensional simplex analysis can be more powerful than node-only approaches, even in cases where the underlying graph may not have many higher-dimensional structures.
Diffusion+SVR HGLET+SVR GHWT+SVR SchNet PaiNN SO3Net I SO3Net II
@@
Table 5: Comparison of the performance of our MHSNs and the other state-of-the-art GNNs for potential energy prediction. We report the accuracy via MAE and RMSE as well as the number of trainable parameters in each network.